# Learnability in Online Kernel Selection with Memory Constraint via Data-dependent Regret Analysis

### v1

@@ 3.1 Adaptive Sampling
We use the optimistic mirror descent framework (OMD) [8, 9] to generate $\{f_{t,i}\in\mathcal{H}_{i}\}^{T}_{t=1}$. OMD maintains two sequences of hypothesis, i.e., $\{f_{t,i}\}^{T}_{t=1}$ and $\{f^{\prime}_{t,i}\}^{T}_{t=1}$ which are defined as follows,
@@
where $\nabla_{t,i}$ is the (sub)-gradient, ^$\hat{\nabla}_{t,i}$ is an optimistic estimator of $\nabla_{t,i}$ and
@@
Note that the hypothesis space in (4) and (5) is $\mathcal{H}_{i}$ and $\mathbb{H}_{i}$, respectively. This trick [23] is critical to obtaining regret bounds depending on kernel alignment. At the beginning of round $t$, we execute (4) and output $f_{t,i}({\bm{x}}_{t})$. The key is to control the size of $f^{\prime}_{t,i}$. Using Assumption 1, we construct a buffer denoted by $S_{i}$ which stores the examples constructing $f^{\prime}_{t,i}$. Thus we just need to limit the size of $S_{i}$. To this end, we propose an adaptive sampling scheme. Let $b_{t,i}$ be a Bernoulli random variable and
@@
~$\tilde{\nabla}_{t,i}=$ $\frac{\nabla_{t,i}-\hat{\nabla}_{t,i}}{\mathbb{P}[b_{t,i}=1]}\mathbb{I}_{b_{t,i}=1}+\hat{\nabla}_{t,i}.$
@@ 3.2 Removing a Half of Examples
We allocate a memory budget $\mathcal{R}_{i}$ for $f^{\prime}_{t,i}$ and require $\sum^{K}_{i=1}\mathcal{R}_{i}\leq\mathcal{R}$. Using Assumption 1, we just need to limit $|S_{i}|\leq\alpha\mathcal{R}_{i}$. At any round $t$, if $|S_{i}|=\alpha\mathcal{R}_{i}$ and $b_{t,i}=1$, then we remove a half of examples from $S_{i}$ [23]. Let $S_{i}=\{({\bm{x}}_{r_{j}},y_{r_{j}})\}^{\alpha\mathcal{R}_{i}}_{j=1}$. We rewrite $f^{\prime}_{t-1,i}$ as follows,
@@
$f^{\prime}_{t-1,i}(1)=$ $\sum^{\alpha\mathcal{R}_{i}/2}_{j=1}\beta_{r_{j}}\kappa({\bm{x}}_{r_{j}},\cdot),$
$f^{\prime}_{t-1,i}(2)=$ $\sum^{\alpha\mathcal{R}_{i}}_{j=\alpha\mathcal{R}_{i}/2+1}\beta_{r_{j}}\kappa({\bm{x}}_{r_{j}},\cdot).$
@@ 3.3 Reduction to PEA
Similar to previous online multi-kernel learning algorithms [1, 10, 24], we use PEA framework to aggregate the outputs of $K$ hypotheses. Let ${\bm{p}}_{t}\in\Delta_{K-1}$ be a probability distribution over $\{f_{t,i}\}^{K}_{i=1}$. At each round $t$, we output the prediction $f_{t}({\bm{x}}_{t})=\sum^{K}_{i=1}p_{t,i}f_{t,i}({\bm{x}}_{t})$ or $\hat{y}_{t}=\mathrm{sign}(f_{t}({\bm{x}}_{t})$. We compute ${\bm{p}}_{t+1}$ using the multiplicative-weight update (Subsection 2.1 in [11]),
@@
where $c_{t,i}=g(f_{t,i}({\bm{x}}_{t}),y_{t})$, $g:\mathbb{R}^{2}\rightarrow\mathbb{R}^{+}\cup\{0\}$ and will be defined later.
@@ 4 Applications
4.1 Hinge Loss Function
We consider online binary classification tasks. Let the loss function be the Hinge loss, implying
$\nabla_{t,i}=-y_{t}\kappa_{i}({\bm{x}}_{t},\cdot)\cdot\mathbb{I}_{y_{t}f_{t,i}({\bm{x}}_{t})<1}.$
We use the “Reservoir Sampling (RS)” technique [14, 25] to create ^$\hat{\nabla}_{t,i}$. To this end, we create a new buffer $V$ with size $M\geq 1$. At the beginning of round $t$, we define
^$\hat{\nabla}_{t,i}=-\frac{1}{|V|}\sum_{({\bm{x}},y)\in V}y\cdot\kappa_{i}({\bm{x}},\cdot).$
RS is defined follows. At any round $t$, we add $({\bm{x}}_{t},y_{t})$ into $V$ with probability $\min\{1,\frac{M}{t}\}$. If $|V|=M$ and we will add the current example, then an old example must be removed from $V$ uniformly. We create another buffer $S_{0}$ that stores all of the examples added into $V$. It is easy to prove that $\mathbb{E}[|S_{0}|]\leq M(1+\ln{T})$. Let $\mathcal{R}_{0}=\frac{M(1+\ln{T})}{\alpha}$ and $\mathcal{R}_{i}=\frac{\mathcal{R}-\mathcal{R}_{0}}{K}$ where we require $\mathcal{R}\geq 2\mathcal{R}_{0}$. It is easy to verify that $\sum^{K}_{i=0}\mathcal{R}_{i}=\mathcal{R}$.
@@ 4.1 Hinge Loss Function
The additional ^$\|\hat{\nabla}_{t,i}\|^{2}_{\mathcal{H}_{i}}$ in $Z_{t}$ makes $\mathbb{P}[b_{t,i}=1]<1$ and is important for controlling the times of removing operation, while it also increases the regret. The trick that uses different hypothesis spaces in (4) and (5), gives a negative term in the regret bound. The negative term can cancel out ^$\|\hat{\nabla}_{t,i}\|^{2}_{\mathcal{H}_{i}}$. Let $o_{t}=2$. We remove $f^{\prime}_{t-1,i}(2)$ from $f_{t,i}$. Then (9) becomes
@@
Let $M>1$ and $\alpha\mathcal{R}:=B\geq 2M(1+\ln{T})$. For any $\mathcal{I}_{T}$, the expected times that M-OMD-H executes removing operation on $S_{i}$ are $\left\lceil\frac{12K}{Bk_{1}}\mathcal{A}_{T,\kappa_{i}}(1+\frac{\ln{T}}{M})\right\rceil$ at most.
Note that that the times of removing operation depend on the kernel alignment, not the time horizon $T$.
@@
Let $\mathcal{R}$ satisfy Lemma 1 and $B>K$. Let $U>1$, $\lambda_{i}=U\frac{\sqrt{K}}{\sqrt{2B}}$, $\eta_{t}$ follow (10) and
@@
$\forall\kappa_{i}\in\mathcal{K},f\in\mathbb{H}_{i},\quad\mathbb{E}\left[\mathrm{Reg}(f)\right]=O\left(\sqrt{UL_{T}(f)\ln{K}}+\frac{U\sqrt{B}}{\sqrt{K}}+\frac{\sqrt{K}U\mathcal{A}_{T,\kappa_{i}}\ln{T}}{\sqrt{B}k_{1}}+U\ln{K}\right).$
For any competitor $f$ satisfying $\|f\|_{\mathcal{H}_{i}}<1$, it must be $L_{T}(f)=O(T)$. The upper bound is trivial. To avoid this issue, we must set $U>1$ and exclude such competitors. Theorem 1 reveals how the regret bound increases with $K$, $\mathcal{R}$ and $\mathcal{A}_{T,\kappa_{i}}$. The larger the memory budget is, the smaller the regret bound will be. Theorem 1 gives an answer to Q1. If the optimal kernel $\kappa^{\ast}\in\mathcal{K}$ matches well with the examples, i.e., $\mathcal{A}_{T,\kappa^{\ast}}=o(T)$, then it is possible to achieve a $o(T)$ regret bound in the case of $\mathcal{R}=\Theta(\ln{T})$. Such a result gives an answer to Q2. According to Definition 2, learning is possible in $\mathbb{H}_{\kappa^{\ast}}$. The information-theoretical cost that the optimal kernel is unknown is $O(\sqrt{UL_{T}(f)\ln{K}})$, which is a lower order term.
Our regret bound can also derive the state-of-the-art results. Let $B=\Theta(\mathcal{A}_{T,\kappa_{i}})$. Then we obtain a $O(U\sqrt{K\mathcal{A}_{T,\kappa_{i}}})$ expected bound. Previous work proved a $O(U\sqrt{K}T^{\frac{1}{4}}\mathcal{A}^{\frac{1}{4}}_{T,\kappa_{i}})$ high-probability bound [3]. If $\mathcal{A}_{T,\kappa_{i}}=O(T)$, then we obtain a $O\left(\frac{U\sqrt{K}T}{\sqrt{B}}\right)$ expected bound which matches the upper bound in [4]. If $B=\Theta(T)$, then our regret bound matches the upper bounds in [1, 2, 10].
Next we show an algorithm-dependent bound which proves that removing a half of examples can be better than the restart technique. Let $\mathcal{J}_{i}=\{t\in[T]:|S_{i}|=\alpha\mathcal{R},b_{t,i}=1\}$ and
@@
There must be a constant $\xi_{i}\in(0,4]$ such that $\Lambda_{i}=\xi_{i}U^{2}|\mathcal{J}_{i}|$. Recalling that $\bar{f}^{\prime}_{t-1,i}(1)$ is the initial hypothesis after a removing operation. It is natural that if $(\bar{f}^{\prime}_{t-1,i}(1)-f)$ is close to $(f^{\prime}_{t-1,i}-f)$, then $\xi_{i}\ll 4$. The restart technique sets $\bar{f}^{\prime}_{t-1,i}(1)=0$ which implies $\Lambda_{i}\leq U^{2}|\mathcal{J}_{i}|$. In the worst case, our approach is slightly worse than restart. If $\xi_{i}$ is small enough, then our approach is much better than restart.
@@
Let the conditions in Theorem 1 be satisfied. For each $i\in[K]$, let $\lambda_{i}=\frac{\sqrt{2}U}{\sqrt{5\tilde{\mathcal{A}}_{T,\kappa_{i}}}}$. If $\xi_{i}\cdot|\mathcal{J}_{i}|\leq 1$, then the expected regret of M-OMD-H satisfies,
$\forall\kappa_{i}\in\mathcal{K},f\in\mathbb{H}_{i},\quad\mathbb{E}\left[\mathrm{Reg}(f)\right]=O\left(\sqrt{L_{T}(f)\ln{K}}+\frac{UK}{B}\mathcal{A}_{T,\kappa_{i}}\ln{T}+U\sqrt{\mathcal{A}_{T,\kappa_{i}}\ln{T}}+U\ln{K}\right).$
The dominate term in Theorem 2 is $\tilde{O}\left(\frac{UK\mathcal{A}_{T,\kappa_{i}}}{B}\right)$, while the dominate term in Theorem 1 is $\tilde{O}\left(\frac{UK\mathcal{A}_{T,\kappa_{i}}}{\sqrt{B}}\right)$. The restart technique can only enjoys a regret bound given in Theorem 1. Thus removing a half of examples can be better than restart.
@@ 4.2 Smooth Loss Functions
Assumption 3.
[19] Let $G_{1}$ and $G_{2}$ be positive constants. For any $u$ and $y$, $\ell(u,y)$ satisfies (i) $|\ell^{\prime}(u,y)|\leq G_{1}$, (ii) $|\ell^{\prime}(u,y)|\leq G_{2}\ell(u,y)$, where $\ell^{\prime}(u,y)=\frac{\mathrm{d}\,\ell(u,y)}{\mathrm{d}\,u}$.
The logistic loss function satisfies Assumption 3 with $G_{2}=1$. For the hinge loss function, M-OMD-H uniformly allocates the memory budget and induces a $O(\sqrt{K})$ factor in the regret bound. For smooth loss functions, we will propose a memory sharing scheme that can avoid the $O(\sqrt{K})$ factor.
Recalling that the memory is used to store the examples and coefficients. The main challenge is how to share the examples. To this end, we only keep a single buffer for $\{f_{t,i}\}^{K}_{i=1}$. We first rewrite $f_{t}({\bm{x}}_{t})$ as follows,
@@
Thus we just need to ensure $S_{i}=S$ for all $i\in[K]$. For each $f_{t,i}$, we define a surrogate gradient $\nabla_{t,i}=\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\kappa_{i}({\bm{x}}_{t},\cdot)$. Note that we use the first derivative $\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})$, but not $\ell^{\prime}(f_{t,i}({\bm{x}}_{t}),y_{t})$. Let ^$\hat{\nabla}_{t,i}=0$. We can merge (4) and (5) and give a single update rule
@@
$f_{t+1,i}=\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\langle f,\nabla_{i(s_{t}),i}\rangle+\mathcal{B}_{\psi_{i}}(f,f_{t,i})\right\},$ (13)
where $\nabla_{i(s_{t}),i}=\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\kappa_{i}({\bm{x}}_{i(s_{t})},\cdot)$. Otherwise, we replace (8) with (14).
@@
$c_{t,i}=\left\{\begin{array}[]{ll}\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\cdot\left(f_{t,i}({\bm{x}}_{t})-\min_{j\in[K]}f_{t,j}({\bm{x}}_{t})\right)&~{}\mathrm{if}~{}\ell^{\prime}(f_{t})>0,\\ \ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\cdot\left(f_{t,i}({\bm{x}}_{t})-\max_{j\in[K]}f_{t,j}({\bm{x}}_{t})\right)&~{}\mathrm{otherwise}.\end{array}\right.$ (16)
@@
Let $\delta\in(0,1)$, $\alpha\mathcal{R}:=B\geq 21\ln\frac{1}{\delta}$ and $\ell$ satisfy Assumption 3. For any $\mathcal{I}_{T}$, with probability at least $1-\frac{T}{B}\Theta(\lceil\ln{T}\rceil)\delta$, the times that M-OMD-S executes removing operation on $S$ are $\left\lceil\frac{4G_{2}\hat{L}_{1:T}}{(B-\frac{4}{3}\ln\frac{1}{\delta})G_{1}}\right\rceil$ at most, in which $\hat{L}_{1:T}=\sum^{T}_{t=1}\ell(f_{t}({\bm{x}}_{t}),y_{t})$.
It is important that the times of removing operation depend on the cumulative losses of the algorithm, not the time horizon $T$.
@@
Let $K\leq d$, $\eta_{t}$ follow (10),

### v2

@@ 4.1 Adaptive Sampling
For all $i\in[K]$, we use the optimistic mirror descent framework (OMD) [Chiang2012Online, Rakhlin2013Online] to generate $\{f_{t,i}\}^{T}_{t=1}$. OMD maintains two sequences of hypothesis, i.e., $\{f_{t,i}\}^{T}_{t=1}$ and $\{f^{\prime}_{t,i}\}^{T}_{t=1}$ which are defined as follows,
@@
where $\nabla_{t,i}$ is the (sub)-gradient of $\ell(f_{t,i}({\bm{x}}_{t}),y_{t})$, ^$\hat{\nabla}_{t,i}$ is an optimistic estimator of $\nabla_{t,i}$ and
@@
Note that the hypothesis space in (4) and (5) is $\mathcal{H}_{i}$ and $\mathbb{H}_{i}$, respectively. This trick [Li2023Improved] is critical to obtaining regret bounds depending on kernel alignment. At the beginning of round $t$, we execute (4) and compute $f_{t,i}({\bm{x}}_{t})$. The key is to control the size of $f^{\prime}_{t,i}$. By Assumption 1, we construct a buffer denoted by $S_{i}$ that stores the examples constructing $f^{\prime}_{t,i}$. Thus we just need to limit the size of $S_{i}$. To this end, we propose an adaptive sampling scheme. Let $b_{t,i}$ be a Bernoulli random variable satisfying
@@
~$\tilde{\nabla}_{t,i}=$ $\frac{\nabla_{t,i}-\hat{\nabla}_{t,i}}{\mathbb{P}[b_{t,i}=1]}\cdot\mathbb{I}_{b_{t,i}=1}+\hat{\nabla}_{t,i}.$
@@ 4.2 Removing a Half of Examples
We allocate a memory budget $\mathcal{R}_{i}$ for $f^{\prime}_{t,i}$ and require $\sum^{K}_{i=1}\mathcal{R}_{i}\leq\mathcal{R}$. By Assumption 1, we just need to limit $|S_{i}|\leq\alpha\mathcal{R}_{i}$. At any round $t$, if $|S_{i}|=\alpha\mathcal{R}_{i}$ and $b_{t,i}=1$, then we remove a half of examples from $S_{i}$ [Li2023Improved]. Let $S_{i}=\{({\bm{x}}_{r_{j}},y_{r_{j}})\}^{\alpha\mathcal{R}_{i}}_{j=1}$. We rewrite $f^{\prime}_{t-1,i}$ as follows,
@@
$f^{\prime}_{t-1,i}(1)=$ $\sum^{\alpha\mathcal{R}_{i}/2}_{j=1}\beta_{r_{j}}\kappa_{i}({\bm{x}}_{r_{j}},\cdot),$
$f^{\prime}_{t-1,i}(2)=$ $\sum^{\alpha\mathcal{R}_{i}}_{j=\alpha\mathcal{R}_{i}/2+1}\beta_{r_{j}}\kappa_{i}({\bm{x}}_{r_{j}},\cdot).$
@@ 4.3 Reduction to PEA
Similar to previous online multi-kernel learning algorithms [Sahoo2014Online, Shen2019Random, Jin2010Online], we use PEA framework [Cesa-Bianchi2006Prediction] to aggregate $K$ predictions. Let $\Delta_{K-1}$ be the $K-1$ dimensional simplex. At each round $t$, we maintain a probability distribution ${\bm{p}}_{t}\in\Delta_{K-1}$ over $\{f_{t,i}\}^{K}_{i=1}$. and output the prediction $f_{t}({\bm{x}}_{t})=\sum^{K}_{i=1}p_{t,i}f_{t,i}({\bm{x}}_{t})$ or $\hat{y}_{t}=\mathrm{sign}(f_{t}({\bm{x}}_{t})$. By the multiplicative-weight update in Subsection 2.1 [Cesa-Bianchi2006Prediction], ${\bm{p}}_{t+1}$ is defined as follows
@@
where $c_{t,i}=g(f_{t,i}({\bm{x}}_{t}),y_{t})$. $g(\cdot,\cdot):\mathbb{R}^{2}\rightarrow\mathbb{R}^{+}\cup\{0\}$ will be defined later.
@@ 5 Applications
5.1 the Hinge Loss Function
We consider online binary classification tasks. Let the loss function be the Hinge loss. Thus $\nabla_{t,i}=-y_{t}\kappa_{i}({\bm{x}}_{t},\cdot)\cdot\mathbb{I}_{y_{t}f_{t,i}({\bm{x}}_{t})<1}$. We use the “Reservoir Sampling (RS)” technique [Hazan2009Better, Vitter1985Random] to create ^$\hat{\nabla}_{t,i}$. To this end, we create a new buffer $V_{t}$ with size $M\geq 1$. Initializing ^$\hat{\nabla}_{1,i}=0$, and for $t\geq 2$, we define
^$\hat{\nabla}_{t,i}=-\frac{1}{|V_{t}|}\sum_{({\bm{x}},y)\in V_{t}}y\cdot\kappa_{i}({\bm{x}},\cdot).$
RS is defined follows. At the end of round $t\geq 1$, we add $({\bm{x}}_{t},y_{t})$ into $V_{t}$ with probability $\min\left\{1,\frac{M}{t}\right\}$. If $|V_{t}|=M$ and we will add the current example, then an old example must be removed from $V_{t}$ uniformly. We create another buffer $S_{0}$ that stores all of the examples added into $V_{t}$. It is easy to prove that $\mathbb{E}[|S_{0}|]\leq M(1+\ln{T})$. Let $\mathcal{R}_{0}=\frac{M(1+\ln{T})}{\alpha}$ and $\mathcal{R}_{i}=\frac{\mathcal{R}-\mathcal{R}_{0}}{K}$ where we require $\mathcal{R}\geq 2\mathcal{R}_{0}$. It is easy to verify that $\sum^{K}_{i=0}\mathcal{R}_{i}=\mathcal{R}$.
@@
The additional term ^$\left\|\hat{\nabla}_{t,i}\right\|^{2}_{\mathcal{H}_{i}}$ in $Z_{t}$ is important for controlling the times of removing operation, while it also increases the regret. The trick that uses different hypothesis spaces in (4) and (5), gives a negative term in the regret bound. The negative term can cancel out the increase on the regret. Let $o_{t}=2$. We remove $f^{\prime}_{t-1,i}(2)$ from $f_{t,i}$. Then (9) becomes
@@
Let $M>1$ and $\alpha\mathcal{R}:=B\geq 2M(1+\ln{T})$. For any $\mathcal{I}_{T}$, the expected times that M-OMD-H executes removing operation on $S_{i}$ are $\left\lceil\frac{4K\tilde{\mathcal{A}}_{T,\kappa_{i}}}{Bk_{1}}\right\rceil$ at most, in which
$\tilde{\mathcal{A}}_{T,\kappa_{i}}=1+\sum^{T}_{t=2}\left\|y_{t}\kappa_{i}({\bm{x}}_{t},\cdot)-\frac{\sum_{({\bm{x}},y)\in V_{t}}y\kappa_{i}({\bm{x}},\cdot)}{|V_{t}|}\right\|^{2}_{\mathcal{H}_{i}}.$
Note that that the times of removing operation depend on the kernel alignment, rather than $T$.
@@
Suppose $\mathcal{R}$ satisfies the condition in Lemma 1. Let $B>K$, $U>1$, $\lambda_{i}=U\frac{\sqrt{K}}{\sqrt{2B}}$, $\eta_{t}$ follow (10) and
@@
$c\forall\kappa_{i}\in\mathcal{K},f\in\mathbb{H}_{i},\quad\mathbb{E}\left[\mathrm{Reg}(f)\right]=O\left(\sqrt{UL_{T}(f)\ln{K}}+\frac{U\sqrt{B}}{\sqrt{K}}+\frac{\sqrt{K}U\mathcal{A}_{T,\kappa_{i}}\ln{T}}{\sqrt{B}k_{1}}\right).$
For any competitor $f$ satisfying $\|f\|_{\mathcal{H}_{i}}<1$, it must be $L_{T}(f)=\Theta(T)$, inducing a trivial upper bound. To address this issue, we must set $U>1$ and exclude such competitors. Theorem 1 reveals how the regret bound increases with $K$, $\mathcal{R}$, and $\mathcal{A}_{T,\kappa_{i}}$. The larger the memory budget is, the smaller the regret bound will be. Theorem 1 gives an answer to Q1. If the optimal kernel $\kappa^{\ast}\in\mathcal{K}$ matches well with the examples, i.e., $\mathcal{A}_{T,\kappa^{\ast}}=o(T)$, then it is possible to achieve a $o(T)$ regret bound in the case of $\mathcal{R}=\Theta(\ln{T})$. Such a result gives an answer to Q2. By Definition 2, learning is possible in $\mathbb{H}_{\kappa^{\ast}}$. The information-theoretical cost that the optimal kernel is unknown is $O\left(\sqrt{UL_{T}(f)\ln{K}}\right)$, which is a lower order term.
Our regret bound can also derive the state-of-the-art results. Let $B=\Theta(\mathcal{A}_{T,\kappa_{i}})$. Then we obtain a $O\left(U\sqrt{K\mathcal{A}_{T,\kappa_{i}}}\right)$ expected bound. Previous work proved a $O\left(U\sqrt{K}T^{\frac{1}{4}}\mathcal{A}^{\frac{1}{4}}_{T,\kappa_{i}}\right)$ high-probability bound [Liao2021High]. If $\mathcal{A}_{T,\kappa_{i}}=O(T)$, then we achieve an expected bound of $O\left(U\sqrt{K}\frac{T}{\sqrt{B}}\right)$, which matches the upper bound in [Li2022Worst]. If $B=\Theta(T)$, then our regret bound matches the upper bounds in [Sahoo2014Online, Foster2017Parameter, Shen2019Random].
Next, we present an algorithm-dependent bound demonstrating that removing half of the examples can be more effective than the restart technique. We introduction two new notations as follows
$\mathcal{J}_{i}=$$\{t\in[T]:|S_{i}|=\alpha\mathcal{R},b_{t,i}=1\},$
@@
There must be a constant $\xi_{i}\in(0,4]$ such that $\Lambda_{i}=\xi_{i}U^{2}|\mathcal{J}_{i}|$. Recalling that $\bar{f}^{\prime}_{t-1,i}(1)$ is the initial hypothesis after a removing operation. It is natural that if $\left(\bar{f}^{\prime}_{t-1,i}(1)-f\right)$ is close to $\left(f^{\prime}_{t-1,i}-f\right)$, then $\xi_{i}\ll 4$. The restart technique sets $\bar{f}^{\prime}_{t-1,i}(1)=0$ implying $\Lambda_{i}\leq U^{2}|\mathcal{J}_{i}|$. In the worst case, our approach is slightly worse than the restart. If $\xi_{i}$ is sufficiently small, then our approach is much better than the restart.
@@
Suppose the conditions in Theorem 1 are satisfied. For each $i\in[K]$, let $\lambda_{i}=\frac{\sqrt{2}U}{\sqrt{5\tilde{\mathcal{A}}_{T,\kappa_{i}}}}$. If $\xi_{i}\cdot|\mathcal{J}_{i}|\leq 1$ for all $i\in[K]$, then the expected regret of M-OMD-H satisfies,
$\forall\kappa_{i}\in\mathcal{K},f\in\mathbb{H}_{i},\quad\mathbb{E}\left[\mathrm{Reg}(f)\right]=O\left(\sqrt{L_{T}(f)\ln{K}}+\frac{UK}{B}\mathcal{A}_{T,\kappa_{i}}\ln{T}+U\sqrt{\mathcal{A}_{T,\kappa_{i}}\ln{T}}\right).$
The dominate term in Theorem 2 is $\tilde{O}\left(\frac{UK\mathcal{A}_{T,\kappa_{i}}}{B}\right)$, while the dominate term in Theorem 1 is $\tilde{O}\left(\frac{UK\mathcal{A}_{T,\kappa_{i}}}{\sqrt{B}}\right)$. The restart technique only enjoys the regret bound given in Theorem 1. Thus removing a half of examples can be better than the restart.
@@ 5.2 Smooth Loss Functions
Assumption 3 ([Zhang2013Online]).
Let $G_{1}$ and $G_{2}$ be positive constants. For any $u$ and $y$, $\ell(u,y)$ satisfies (i) $|\ell^{\prime}(u,y)|\leq G_{1}$, (ii) $|\ell^{\prime}(u,y)|\leq G_{2}\ell(u,y)$, where $\ell^{\prime}(u,y)=\frac{\mathrm{d}\,\ell(u,y)}{\mathrm{d}\,u}$.
The logistic loss function satisfies Assumption 3 with $G_{2}=1$. For the hinge loss function, M-OMD-H uniformly allocates the memory budget and induces a $O\left(\sqrt{K}\right)$ factor in the regret bound. For smooth loss functions, we will propose a memory sharing scheme that can avoid the $O\left(\sqrt{K}\right)$ factor.
Recalling that the memory is used to store the examples and coefficients. The main challenge is how to share the examples. To this end, we only keep a single buffer $S$. We first rewrite $f_{t}({\bm{x}}_{t})$ as follows,
@@
We just need to ensure $S_{i}=S$ for all $i\in[K]$. For each $f_{t,i}$, we define a surrogate gradient $\nabla_{t,i}=\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\cdot\kappa_{i}({\bm{x}}_{t},\cdot)$. Note that we use the first derivative $\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})$, rather than $\ell^{\prime}(f_{t,i}({\bm{x}}_{t}),y_{t})$. Let ^$\hat{\nabla}_{t,i}=0$ in (4). Then we obtain a single update rule as follows
@@
$\begin{split}f_{t+1,i}=&\mathop{\arg\min}_{f\in\mathbb{H}_{i}}\left\{\left\langle f,\nabla_{i(s_{t}),i}\right\rangle+\mathcal{B}_{\psi_{i}}(f,f_{t,i})\right\},\\ \nabla_{i(s_{t}),i}=&\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\cdot\kappa_{i}\left({\bm{x}}_{i(s_{t})},\cdot\right).\end{split}$ (13)
Otherwise, we replace (8) with (14).
@@
$c_{t,i}=\left\{\begin{array}[]{ll}\ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\cdot\left(f_{t,i}({\bm{x}}_{t})-\min_{j\in[K]}f_{t,j}({\bm{x}}_{t})\right),&~{}\mathrm{if}~{}\ell^{\prime}(f_{t})>0,\\ \ell^{\prime}(f_{t}({\bm{x}}_{t}),y_{t})\cdot\left(f_{t,i}({\bm{x}}_{t})-\max_{j\in[K]}f_{t,j}({\bm{x}}_{t})\right),&~{}\mathrm{otherwise}.\end{array}\right.$ (16)
@@
Suppose $\ell(\cdot,\cdot)$ satisfies Assumption 3. Let $\delta\in(0,1)$ and $\alpha\mathcal{R}:=B\geq 21\ln\frac{1}{\delta}$. For any $\mathcal{I}_{T}$, with probability at least $1-\frac{T}{B}\Theta(\lceil\ln{T}\rceil)\delta$, the times that M-OMD-S executes removing operation on $S$ are $\left\lceil\frac{4G_{2}\hat{L}_{1:T}}{(B-\frac{4}{3}\ln\frac{1}{\delta})G_{1}}\right\rceil$ at most, in which
$\hat{L}_{1:T}=\sum^{T}_{t=1}\ell(f_{t}({\bm{x}}_{t}),y_{t}).$
It is important that the times of removing operation depend on the cumulative losses of the algorithm, rather than $T$.
@@
Suppose $\mathcal{R}$ and $\ell(\cdot,\cdot)$ satisfy Lemma 2. Let $\delta\in(0,1)$, $K\leq d$, $\eta_{t}$ follow (10),